OpenStudy (anonymous):

Bag A contains 3 green, 2 blue and 5 red balls and bag B contains 2 green, 3 red and 4 blue balls. A dice is rolled, if it gives a multiple of 3, bag A will be selected and 2 balls will be drawn. If the dice gives a multiple of 5, bag B will be selected and 2 balls will be drawn. If the dice gives neither a multiple of 5 nor a multiple of 3, one ball is drawn from each bag. What is the probability that both the balls are blue?

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

2/171

OpenStudy (anonymous):

srry i couldnt able to delete the above one

OpenStudy (anonymous):

my delete button is not working

OpenStudy (kropot72):

What are the three values of probability for the three possible outcomes of rolling the dice?

OpenStudy (anonymous):

2 blue balls from bag a and 4 blue balls in bag b so we should pick two

OpenStudy (anonymous):

and also 2 balls will be drawn from bag a when it is a multiple of 3 and 2 balls from bag b will be drawn so multiple of 5

OpenStudy (kropot72):

You need to answer my question to enable me to help you. "What are the three values of probability for the three possible outcomes of rolling the dice?" 1. What is the probability of rolling a multiple of 3? 2. What is the probability of rolling a multiple of 5? 3 What is the probability of rolling neither a multiple of 5 nor a multiple of 3?

OpenStudy (anonymous):

1)2/6 2)1/6 3)4/6

OpenStudy (anonymous):

1/3,1/6,2/3

OpenStudy (kropot72):

What is the probability of rolling neither a multiple of 5 nor a multiple of 3? There are 3 outcomes that are neither a multiple of 5 nor a multiple of 3. They are 1,2 and 4. So what is the correct value of probability?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

oh srry i included 6 also

OpenStudy (anonymous):

1/3*1/171+1/6*6/171

OpenStudy (kropot72):

Consider the first case: P(multiple of 3) = 1/3 If a multiple of 3 is rolled, what is the probability of drawing two blue balls from bag A? There are initially 10 balls in bag A. The probability of drawing a blue ball on the first draw is 2/10. If a blue ball is drawn first, the probability of drawing a blue ball on the second draw is 1/9. The probability of rolling a multiple of 3 and drawing 2 blue balls from bag A is given by: \[\large \frac{1}{3}\times\frac{2}{10}\times\frac{1}{9}=\frac{2}{270}\] Do you understand this step?

OpenStudy (anonymous):

yes.1/135

OpenStudy (anonymous):

from bag b means 1/6*4/9*3/8=12/432

OpenStudy (anonymous):

1/36

OpenStudy (anonymous):

so 4+15/540=19/540

OpenStudy (kropot72):

You are correct for the second case! Can you find the probability of the third case?

OpenStudy (anonymous):

i think we dont need third case cuz if neither 5 or 3 so only one ball is drawn and not from a or b bag.we need two balls that to blue balls

OpenStudy (anonymous):

19/540

OpenStudy (kropot72):

Yes the third case is needed. Two balls are drawn, one ball from each bag A and bag B.

OpenStudy (anonymous):

1/2*2/10*4/9

OpenStudy (kropot72):

The probability of the third case is given by: \[\large \frac{1}{2}\times\frac{2}{10}\times\frac{4}{9}=\frac{2}{45}\] The three cases are mutually exclusive. Therefore the probability of 2 blue balls is: \[\large P(2\ blue\ balls)=\frac{1}{135}+\frac{1}{36}+\frac{2}{45}=you\ can\ calculate\]

OpenStudy (anonymous):

(20+75+120)/2700

OpenStudy (anonymous):

215/2700

OpenStudy (anonymous):

43/540

OpenStudy (anonymous):

mutually exclusive means

OpenStudy (kropot72):

\[\large \frac{1}{135}+\frac{1}{36}+\frac{2}{45}=\frac{4+15+24}{540}=\frac{43}{540}\] You are correct!

OpenStudy (kropot72):

Events are 'mutually exclusive' if it is impossible for them to happen together. In your question the three cases must happen separately.